OpenStudy (anonymous):

I have to solve this equation involving rational expressions. Then I have to identify each equation as an identity, an inconsistent equation, or a conditional equation. My first step is to find the LCD. My problem is that I don't know how to with this equation. Can anyone help me find the LCD and explain to me how they got it? :) 1/5x-1/4x+1/3x=(-17)/60

OpenStudy (e.mccormick):

So \(\dfrac{1}{5x}-\dfrac{1}{4x}+\dfrac{1}{3x}=\dfrac{-17}{60}\)?

OpenStudy (anonymous):

Yes

OpenStudy (cwrw238):

the denominators are 5x, 4x, 3x and 60 the LCD is the lowest common denominator into which each of the 4 values will divide. we obviously need to have x in the LCD and the biggest number is 60 into which 3 4 and 5 will all divide so our LCD is 60x

OpenStudy (e.mccormick):

OK. Well, if you look at the denominators, there is a constant and an x. So the x can be seen as a factor and the constant as a factor.

OpenStudy (anonymous):

Okay, thank you very much to the both of you. I understand now. :) I really do appreciate it.

OpenStudy (e.mccormick):

np. Have fun!